OpenStudy (anonymous):

Write algebraic expression for cos(sin-1(x))

OpenStudy (anonymous):

 so cos(sin^-1x) = \[\frac{ \sqrt{1-x^2} }{ 1 }= \sqrt{1-x^2}\]